MODS2 Commissioning Night 4 - 20140503 UT

Observers: D. Miller (LBTO), R. Pogge (OSU), O. Kuhn (LBTO), J. Hill, M. Wagner (LBTO, Remote from Tucson)
Telescope Operator: S. Allanson (LBTO)
AO Support: D. Miller (LBTO)
Telescope Support: John Little et al. (LBTO)
Software Support: T. Golota (Remote from Tucson)
Instrument Support: J. Power (LBTO)

Summary

Active Optics Status

Active Optics is generally working on MODS2 both on-axis and off-axis, but there seems to be four problems to be solved to get active optics really running smoothly with the MODS2 wfs:
  1. These zernike calibrations are showing that the gains are different by a factor of 1.4 - 2 compared to the other AGWs.
  2. sextractor is not doing a good job of measuring centroids of the spots. This results in lots of zernike noise.
  3. GCS needs to run this trimmed 9x9 mask that IDL is running to avoid the spots on the diagonals that are on the edge of the mirror.
  4. There may be an SVD decomposition problem where there is too much redundancy between Z4,Z11,Z22.

Details

Conditions at sunset: clear skies, a few wispy clouds low on the W and S horizons.

Collimation Test in Bright Twilight

~02:15 Open and "collimating" on a bright star. First twilight sky WFS image is /Repository/20140503/GCS/right_wfscimage000001.fits
  • First WFSC images, which are off-axis (x=0,y=-162.5) and in twilght:
    20140503_firstWFSCimages_offaxis.jpg
The left 3 columns of the wfs appear to be vignetted.

02:34 Restart GCSR

On-Axis Active Optics

02:40 Collimating on BS9142 on-axis.
  • right_wfscimage000014 on...
    first_onaxis.jpg

02:46 Changing wfs x from 417.6 to 430 messes up the wfs cutout image. Maybe we needed to restart GCS for that to be correct?

02:53 Doug measured new pupil center x=408.0 y=222.0.

03:00 Collimation ran-away. John and Doug discussing both missing columns of spots on the left and scale factors.
  • Rick noting that the WFSCam_pixelscale in MODS_R.cfg should be different from that used in MODS_L.cfg -smaller by a factor 25/30. John changed the value from 0.223881 to 0.1866 the following night.

Doug and John adjusting lots of things to improve collimation. (Literally fixing half a dozen minor issues and parameter changes)

03:16 Use PSFGUI to reduce Active Optics Gains on Z5,Z6,Z9,Z10 to suppress the apparent oscillation.

03:28 Reset the AdSec flat and clear active optics. SH spots look a bit better.

03:37 Z5 and Z6 Active Optics gain to 0.1

Target is BS9142

03:43 John changed rotAngleDir from +1 to -1 as the comment suggests, but it didn't do anything. He put it back to normal after 10-15 minutes.

03:45 Clear M1 forces, Z11, Z22 Active Optics gains to 0, Z5, Z6 gain to 0.2

03:48 Clear M2 active optics - The problem here seems to be that Active Optics is driving us to a non-zero converged amount of coma.

03:54 Send +500 nm Z7 to M2, and active optics removes it in the first step (saying -1021 nm).

03:57 Send +500 nm Z8 to M2, and it also is driven out.

04:00 Send +2000 nm Z4 which was a bit too small.

Off-Axis Guiding and Active Optics

04:05 ACTIVE preset off-axis x=0, y=-87. Pupil moved either the wrong direction or too far in the cutout, and appears to be off by a full column. See wfsc 0147.

04:08 ACTIVE preset on-axis. See wfsc 0162.

Doug says the lenslet spacing is not the usual 13.38 pixels, but rather 13.2 pixels. This is one of the multiple small things that we have fixed on MODS_R.cfg.

04:19 ACTIVE preset changing PA by 90 deg. See wfsc 0166

04:23 ACTIVE preset back to the previous default PA with 13.2 pixels per subap.

Adjusting the Guidecam Hotspot

04:35 Adjusting hotspot using setHotspot right, but GCSR is not using the modified value. We think Taras broke this in the current or an earlier build. See Issue 5168.

04:39 All the Active Optics Gains in PSF are back to 1.0.

04:43 ACTIVE preset to get parameter updates from config file. It updated guidecam_center_x/y, but did not update guidecam_hotspot_x/y. So we have to restart GCS on every adjustment. Prior to this point, hotspot was always 269,234.

04:47 hotspot 290,240 - see wfsc 0203 (too far up on wfs)

Moving the hotspot in X adjust the star on the wfs in Y.

04:54 hotspot 265, 215 (too far left on wfs)

04:57 hotspot 265, 255 (too far right on wfs)

05:00 hotpsot 265, 245 (slightly low, slightly left)

05:04 hotspot 267,240 is the adopted value in MODS_R.cfg

Adjusting the Pupil Shift Parameters

05:09 ACTIVE preset off-axis x=0,y=-87

05:14 ACTIVE preset gs=9 x=0, y=-147

Doug applies a smaller pupil shift parameter.

05:28 ACTIVE preset off-axis x=0,y=-87

05:36 ACTIVE preset with PA shifted by 30 deg, x=+43,y=-75

05:40 ACTIVE preset PA=+115, x=75, y=-43

05:44 ACTIVE preset PA=-5, x=-75, y=-43

05:49 ACTIVE preset PA=137.3, but that is too far. -180 mm in Y is the bottom limit.

05:50 ACTIVE preset PA=113.7, x=0 y=-179 - no evidence of vignetting here.

05:56 ACTIVE preset PA=143.7, x=-89.2 y=-155 --> Why are the upper left SH spots elongated? Also note that we are driving in Z11 at the level of 1000 nm.

06:11 ACTIVE preset PA=83.7, x=89.6 y=-154.8

06:20 ACTIVE preset on-axis -- most of the odd spot elongations are gone. How are the off-axis elongations created by the wfs? The Adaptive Secondary has not moved.

Transform Data

06:30 Stone_K3
  • Starting collection of transform data.
    • cd /home/lbto/data/20140503/transform_k3
    • iif_register, /tcs
    • iif_preset, /right, mode=active, Stone_K3,gs=0
    • transform_collect_new
      • red dual IMCS is not locking at PA=90.

07:15 Stone_L3 (more stars and PA=0,90,180,270 all available).
  • iif_preset,mode=active,Stone_L3,gs=0,pa=0
  • converged & IMCS locked, seeing ~ 0.85"
  • stop guiding and pull out guide probe
    • 3Kx3K: mods2r.20140503.0013.fits (10-sec is good)
    • 1Kx1K: mods2r.20140503.0014.fits (10-sec)
  • transform_collect_new ---> select Stone_L3_gs23.list
    • starting data collection... 07:14 UT (mods2r.20140503.0015.fits ...)
      • note - image 16 is offset 30 0 from 15
      • 17 on is remainder of sequence and there are no offsets here.

07:50 Preset on axis with PA=180 in preparation for next set (Stone_L3_gs9_n5.list)
  • transform_collect_new ---> select Stone_L3_gs9_n5.list (mods2r.20140503.0030.fits)
    • starting data collection... 07:55 UT (mods2r.20140503.0031.fits ...)
08:20 Next field: exit, restart IDL to take the header change
  • iif_register,/tcs
  • iif_preset, /right, mode=active, Stone_L3,gs=30,ts=30,pa=90
    • imcs failed to lock here... giving up on PA=90
  • iif_preset, /right, mode=active, Stone_L3, gs=0,ts=0,pa=0 (no --- want pa=270)
  • iif_preset, /right, mode=active, Stone_L3, gs=0, ts=0, pa=270
    • imcslock, wfs collimated, pull out guide probe. take image. (mods2r.20140503.0039.fits)
  • transform_collect_new ---> select Stone_L3_n6.list (mods2r.20140503.0040.fits on ...)

08:49 Completion of Transform Data

7x7 Mask on Active Optics

08:50 Stop and restart GCSR

08:51 ACTIVE preset on-axis - "missing Reconstruction matrix"

Spent a while sorting this out and restarted GCS a few times.

09:04 ACTIVE preset works fine after switching back to the old 9x9 configuration.

Guider Astrometric Calibration

09:15 ACTIVE preset to M92 17:17:05.2 43:08:59.31 to check pointing

ACQUIRE preset to star near field

ACTIVE preset on nearby star to collimate

GUIDE preset on-axis on field (this may have changed pointing because of density of field)
  • stop guiding
  • take data:
    • Rick is using readGuideCam -e t(ms) to take images and is changing filters via the MODS GUI.
      • these are right_guiderimage.0000NN NN=3-22 and in /home/modseng/20140503/ (later copied to /home/lbto/MODS2Comm/GuiAstrom/Data/20140503/)
      • 10 LP filter 5-sec
      • 11-12 B-Bessel 5-10 sec
      • 13-14 clear 5-sec
      • 15 ND1.0 filters
      • 16-18 ND 1.0 10-20 sec
    • Rick took a series of biases and flats (closed dome with dome lights on)
    • Adding more detailed notes, finding charts to GuideCamAstPhot

09:40 Done with Guider astrometry. John and Doug working on Zernikes now.

Zernike Calibration

On BS9169 on-axis

data in /home/lbto/data/20140503/zernike/

09:41 and 09:45 Two ACTIVE presets 180 deg apart in PA show that there is a bit of pupil wobble.

stopWFS

09:50 zernike_collect with sext_4_22.list - there is a problem with sextractor spots in IDL not matching the eyeball centroids of the wfsc image.

10:04 Doug is running zernike_collect with 9x9 mask

10:12 Now a better mask which is 9x9 with the corners cut off. Still a lot of oscilation of trefoil.

10:17 ACTIVE preset with GCS on-axis - Z9,Z10 Active Optics Gains turned down to 0.2 on PSF.

Doug changed a couple sextractor parameters.

10:25 zernike_collect - it isn't converging below 600 nm rms.

10:31 zernike_collect, limit=800, /verbose with all_sext.list (which has larger amplitudes)

Sext
Zernike Amplitude Measured Average Factor wfsc #
0 0 110     1-2
4 1000.00 1235 1035 996 1088 0.92 3-5
0 0 -604     6
0 0 -76     6-10
5 1000.00 1160 1076 1227 1155 0.87 11-13
0 0 -245     14
0 0 43     14
6 1000.00 1236 1328 1332 1299 0.77 15-17
0 0 -343     18
0 0 -164     18-21
7 1000.00 -984 -714 -1008 -902 -1.11 22-24
0 0 186     25
0 0 -202     25-27
8 1000.00 -1298 -1153 -1438 -1296 -0.77 28-30
0 0 -155     31
0 0 217     31-32
9 1000.00 1423 730 772 975 1.03 33-35

10:54 Primary force adjustment fails during Z10 measurement. Stop script. Clear active optics on M1 and M2.

10:56 zernike_collect, limit=800, /verbose, start=7 with sext_4_22.list

Sext
Zernike Amplitude Measured Average Factor wfsc #
0 0 697     42
0 0 271     42-47
7 500.000 -849 -865 -816 -843 -0.59 48-50
0 0 -278     51
0 0 251     51
8 500.000 -529 -695 -760 -661 -0.76 52-54
0 0 -78     55
0 0 15     55
9 500.000 715 385 754 618 0.81 56-58
0 0 257     59
0 0 104     59-60
10 500.000 272 163 246 227 2.20 61-63
0 0 107     64
0 0 -482     64
11 500.000 -602 -453 -459 -505 -0.99 65-67
0 0 -823     68
0 0 33     68-69
12 300.000 470 426 208 368 0.81 70-72

11:18 Clear Active Optics on M1 between Z12 and Z13. The Z12 measurement may be bad due to a crosstalk runaway? Or the runaway was only on the measurement following Z12?

Note that Z10 above and Z16 below don't seem to be believable results.

0 0 -134     80
0 0 -21     80-86
13 300.000 318 419 231 323 0.93 87-89
0 0 -302     90
0 0 -49     90
14 300.000 682 436 452 523 0.57 91-93
0 0 53     94
0 0 -6     94
15 300.000 434 453 570 486 0.62 95-97
0 0 285     98
0 0 -126     98
16 300.000 -31 -36 143 25 11.88 99-101
0 0 -320     102
0 0 -559     102
17 300.000 -296 -350 -272 -306 -0.98 103-105
0 0 -573     106
0 0 219     106
18 300.000 759 669 747 725 0.41 107-109
0 0 115     110
0 0 10     110
19 300.000 429 523 659 537 0.56 111-113
0 0 11     114
0 0 -63     114
20 300.000 362 560 522 482 0.62 115-117

11:40 Another crosstalk runaway happened between Z20 and Z21. Clear Active Optics and stop zernike_collect.

11:42 zernike_collect, limit=800, /verbose, start=21 with sext_4_22.list

Sext
Zernike Amplitude Measured Average Factor wfsc #
0 0 -81     123
0 0 -257     123-127
21 300.000 452 613 505 523 0.57 128-130
0 0 -207     131
0 0 400     131
22 300.000 1051 1011 935 999 0.30 132-134
0 0 598     135

Repeating failures from above:

11:xx zernike_collect, limit=800, /verbose, start=10 with sext_4_22.list

Sext
Zernike Amplitude Measured Average Factor wfsc #
0 0 -114     136
0 0 -167     136-137
10 500.000 889 354 614 619 0.81 138-140

11:56 zernike_collect, limit=800, /verbose, start=16 with sext_4_22.list

Crosstalk ran away again........so we didn't get Z16.

Close

11:58 Close for impending dawn

Videocon

We were in a 3-way Skype call with Doug for most of the night, but we note that the new Polycoms cut off the call at 8 hours and 30 sec.
Topic attachments
I Attachment Action Size Date Who Comment
20140503_firstWFSCimages_offaxis.jpgjpg 20140503_firstWFSCimages_offaxis.jpg manage 94 K 03 May 2014 - 02:51 OlgaKuhn First WFSC images, which are off-axis and in twilght
first_onaxis.jpgjpg first_onaxis.jpg manage 32 K 03 May 2014 - 02:52 OlgaKuhn  
Topic revision: r9 - 04 May 2014, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback